# Robust Finetuning of Vision-Language-Action Robot Policies via Parameter Merging

### v1

@@ 6.4 RETAIN enables learning multiple skills robustly
Finally, we test whether RETAIN can enable learning multiple skills in sequence, as described in Section˜5.3, and still retain its generalist abilities. We consider learning the two DROID tasks sequentially, first finetuning on plates, and then using this as an initialization to finetune on whiteboard. As outlined in Section˜5.3, RETAIN uses the merged model from the first stage of finetuning as initialization for the second finetuning stage. During evaluation time, we test whether the final policy, after it has sequentially been trained on both tasks, can solve both tasks under ID and OOD evaluations. We compare against co-FT in the sequential learning setting, since it is the strongest baseline at retaining prior knowledge in single-task finetuning (see Fig.˜6). Fig.˜10 shows the performance of the two policies on the two tasks under both ID and OOD settings. We also plot the performance of these two methods in the single-task finetuning setting in dashed lines. These two comparisons serve as the oracle performance ceiling that we expect on these tasks, and is not meant as baselines for the sequential setting. When evaluated on the first task, plates, RETAIN does much better than co-FT in the sequential setting, showing that it is better at retaining its ability to solve the first task even after a second round of finetuning. When evaluated on the second task, whiteboard, RETAIN is also better than co-FT under ID evaluations. RETAIN outperforms co-FT in the sequential setting under all tasks and evaluation types.
@@ 8 Limitations
While we empirically verified that RETAIN works exceptionally well in helping the finetuned model generalize to out-of-distribution variations of the task, we don’t understand the full scope of the reasons why model parameter merging was able to lead to such generalization. This is an interesting area for future work. We have also included some discussion in Section˜A.9 of some hypothesis and why previous work found model parameter merging effective for vision and language tasks. Additionally, RETAIN involves an important hyperparameter, the merginge coefficient, that can be tuned. While we find in our real world experiments that RETAIN is robust to different values of this parameter, slight tuning this parameter is needed. One avenue of future work is determining a good heuristic of how to choose this value.
@@ A.1 DROID Results Details
Fig.˜6 reports performance of two DROID tasks: whiteboard and plates. Here we present the average of the two tasks. The OOD performance shown here is averaged across the val and test scenes.
@@ A.2 LIBERO Results Details
Fig.˜7 reports the average performance of three LIBERO tasks: pot-on-stove, mugs-on-plates, and items-into-box. Here we present the individual performance for all three tasks in Fig.˜12. The OOD performance shown here are averaged only across the two test scene.
@@ A.3 Ablation on Learning Rate and Gradient Steps
Here, we ablate the learning rate and number of gradient steps we take in the task-FT policy in Fig.˜3 to study whether better hyperparameter choices can reduce or resolve overfitting [^1] Following the best practice from Black et al. [7], we always use a learning rate warmup period of $1000$ steps in LIBERO and a consine decay schedule over $30k$ steps to $1/10$ of the peak learning rate. The learning rate we report here are the peak learning rate. . In Fig.˜3, we use learning rate 2.5e52.5e-5. Here in Fig.˜13, we ablate four different learning rates: one greater than the original and two smaller. We evaluate the model at every $100$ gradient steps to also ablate on the number of gradient steps we take. With a larger learning rate, it’s clear that the overfitting issue is more severe, and the performance on all three kinds of evaluations (ID, OOD, Generalist) go down to near zero. With a smaller learning rate, the model still performs well in ID evaluations, and suffers less from forgetting generalist capabilities (measured from Generalist evaluations). This is to be expected because the finetuned model is closer to the pre-trained model with a smaller learning rate. However, note that with a smaller learning rate, the OOD evaluation performance is also worse. We plot the OOD performance achieved using the original learning rate in Fig.˜3 as the dotted orange line in Fig.˜13, and the gap between the dotted and solid orange line shows the performance gap in OOD evaluation when we lower the learning rate, possibly due to underfitting. This new experiment shows that tuning the learning rate and the number of update steps does not solve the problem of overfitting during finetuning. While lowering the learning rate can retain more generalist knowledge, it does not prevent the gradual loss of it. More importantly, lower the learning rate leads to worse OOD evaluation performance.
@@ A.4 Analysis of Finetuning Path in Parameter Space
To understand why model parameter merging helps, we first try to understand here how the model parameter changes during fine-tuning. In particular, we are interested in understanding how linear the finetuning path is in parameter space, since model parameter merging only moves weights in linear paths.
To start out, we check the colinearity of the vectors $\theta_{i+1}-\theta_{i}$ and $\theta_{i}-\theta_{i-1}$, where $\theta_{i}$ is the parameter of the finetuned checkpoint at gradient steps 100i100*i (i.e. we plot the difference vector every $100$ steps during finetuning). We measure the colinearity as the cosine similarity between the two difference vectors. Fig.˜14 shows this for the four different learning rates we ablated in Fig.˜13. Since no values are close to $1$, this shows that the changes in parameter space is highly non-linear. This is expected since the parameter trajectory of deep neural networks is often highly non-linear.
@@
Next, we attempt to more directly visualize the path/direction of the parameters during finetuning by projecting them down to 2D with Principal Component Analysis (PCA). Specifically, we again consider the difference of the weight vectors, $X_{i}=\theta_{i+1}-\theta_{i}$, at every $100$ steps during finetuning. In Fig.˜15, we plot the first two principal components of $X_{i}$ in blue, and label the points $i$. Indeed, we see that for all learning rates, the direction of the parameters is highly non-linear. With small learning rates, the direction oscillates a lot; with larger learning rates, the direction bends in a certain direction. In addition, we plot the two principal components of the parameter-merged model as well in orange. As expected, since the the merged model takes a linear path and the finetuned one does not, the two models end in in very different places in the parameter space. This shows that model merging achieves a different solution than any checkpoints on the finetuning path.
Finally, we analyze changes in all directions of the difference vectors $X_{i}$, instead of just the two principal components as shown in Fig.˜15. We take the difference vector matrix $Y=[X_{1};X_{2};...,X_{n}]$ and compute the singular values of $YY^{T}$ and Fig.˜16. If the difference vectors lie in a linear path, then all difference vectors would point in the same direction and there should only be $1$ non-zero singular value. However, it’s clear from Fig.˜16 that most singular values are non-zero, showing that the path is non-linear in many dimensions. This generalizes the intuition from Fig.˜15 to more dimensions, and shows that model merging indeed achieves a different solution than the finetuning path.
@@
Figure 16: Singular of the parameter difference during finetuning. Each subplot corresponds to a different learning rate.
All these analysis experiments goes to show that the finetuning path is highly non-linear, and therefore model parameter merging actually results in a different solution than the finetuned models.
@@ A.5.4 Evaluation Details
During evaluations, we additionally binarize the policy’s gripper action’s to 0/1 (open/close), as well as clip action magnitudes. The policy’s action horizon is 10 environment steps, and during evaluation, we set the open loop horizon to 8: so, during evaluation, we receive 10 actions from the policy, execute the first 8, and then request a new action chunk. We execute the predicted actions at a control frequency of 15 Hz.
@@ A.5.5 Evaluation Criteria: Whiteboard
For all whiteboard evaluations, we use the criteria specified in 1 to assign partial success. We perform 10 trials per policy evaluation, for both the ID and OOD evals.
@@ A.5.6 Evaluation Criteria: Plates
For all plates evaluations, we use the criteria specified in 2 to assign partial success. We perform 10 trials per policy evaluation, for both the ID and OOD evals.
@@ A.5.7 Generalist Evaluations Details
As detailed in3 and 4, we measure policies’ generalist capabilities by evaluating them on 44 tasks, distributed throughout 9 distinct scenes and 17 different language instructions. Importantly, to ensure fair comparison, we ensure that the initial conditions, camera angle, ligthing, and all other such factors per task are kept the same across the various policies that we evaluate.
Scene Image Task # Trials Randomization Rubric [Uncaptioned image] put the spoon in the dish rack 4 swap spoon and carrot position, 2 evals each 1: pick up spoon; 1.5: move spoon towards dish rack; 2: put spoon in dish rack (anywhere) put carrot in bowl 4 swap spoon and carrot position, 2 evals each 1: pick up carrot; 1.5: move carrot towards bowl; 2: put carrot in bowl put plate in dish rack 2 randomize initial position of the plate in front of the robot 1: pick up plate; 1.5: move plate towards the dish rack; 2: put plate into dish rack (anywhere)" [Uncaptioned image] wipe the table 2 cloth initially on the left and right side of the open area 1: move down towards cloth; 2: perform lateral ""wiping-style"" motion [Uncaptioned image] put the plate on the table 2 plate initially on different dish rack holders (middle and end) 1: moves towards red plate; 1.5: picks up plate; 2: places / drops plate onto the table [Uncaptioned image] clean up the table 2 randomize initial position of paper ball on table 1: picks up paper ball; 1.5 moves paper ball towards brown bin; 2: puts paper ball into bin [Uncaptioned image] close the drawer 4 two top, two bottom drawer 1: moves towards the drawer; 2: closes drawer [Uncaptioned image] put the stapler on the notebook 2 put stapler in higher and lower position on the table 1: picks up stapler; 2: puts stapler on notebook put stapler in the drawer 4 put stapler in higher and lower position on the table, open top and bottom drawer 1: picks up the stapler 2: puts stapler into the drawer [Uncaptioned image] clean the whiteboard 2 initial eraser position on the left and right of whiteboard 1: pick up eraser; 2: perform wiping motions on the whiteboard; 3: erase the full smiley [Uncaptioned image] put the marker in the cup 4 swap initial position of marker and cup, two local modifications each 1: picks up marker; 1.5: rotates arm to put marker in roughly upright position; 2: puts marker in cup
Table 3: DROID Generalist evaluation tasks grouped by scene. Each task contains associated trial counts, randomization details, and evaluation rubrics.
@@
Table 4: DROID Generalist evaluation tasks grouped by scene (cont.). Each task contains associated trial counts, randomization details, and evaluation rubrics.
@@ A.6.1 LIBERO Pretraining
In order to obtain a base-model to serve as the starting point for RETAIN in LIBERO, we pre-train $\pi_{0}$ on a mixture of LIBERO datasets. Specifically, we use 90 tasks from LIBERO-90, 9 tasks from LIBERO-object, 9 tasks from LIBERO-spatial, and 9 tasks from LIBERO-goal, for a total of 117 tasks in our pretraining dataset. For all LIBERO datasets, pretraining and finetuning, we utilize pre-processed RLDS versions gathered from here and here. These datasets consist of LIBERO demonstrations that have been preprocessed to upscale images, filter out transitions with idle actions, and remove failure trajectories. More details for the pretraining stage itself can be found in A.7. All subsequent training and evaluation uses the normalization stats of this pretraining dataset, applying mean/standard-deviation normalization.
@@ A.6.2 LIBERO Finetuning Datasets
As mentioned in 6.1, we finetune on 3 tasks from LIBERO-10: pot-on-stove, mugs-on-plates, and items-into-basket. For each dataset, we obtain pre-processed and filtered versions as described above. 17 highlights what these finetuning tasks, and thus also our ID evals, look like.
@@ A.6.3 LIBERO Evaluation Details
We use the LIBERO simulator for evaluation. Each ID eval is conducted for 20 episodes. Each OOD eval is conducted with 5 seeds, 10 episodes/seed. 18, 19, and 20 show the 3 types of OOD variations we test for in each of the 3 tasks.
The Generalist evals consists of 20 tasks, 5 each from LIBERO-object, LIBERO-spatial, LIBERO-goal, and LIBERO-90, and each task is tested for 10 episodes. 5 highlights a few tasks per LIBERO eval suite that we use in our generalist evals.
@@ A.7.1 Choosing Merging Coefficient $\alpha$
For each task that we consider (e.g. mugs-on-plate in LIBERO, plates in DROID), we test policy performance on several scenes, which are different variations of the same task with different objects, distractors, and backgrounds etc. To choose what values of $\alpha$ we use, we use one OOD scene as the “validation” scene, and tune the hyperparameter $\alpha$ for best performance on that validation scene. Then, we use the rest of the OOD scenes as the “test” scenes, and report the performance of all methods only on the test scenes. In DROID experiments, we only tune $\alpha\in\{0.25,0.5,0.75\}$ on the validation scenes, while in LIBERO experiments we tune $\alpha\in\{0.9,0.8,0.7,0.6,0.5,0.4,0.3,0.2,0.1\}$ since it is cheap to do so. We find that on DROID $\alpha=0.5$ typically tends to perform well across tasks.
@@ A.7.3 Real-world Experiments’ Hyperparameters
Using the above trained checkpoints, we applied RETAIN, to both task-FT and co-FT, by checking different mergings and recording the best ones. Here are the merging hyperparameters.
@@
Finally, in our continual learning experiment, we use exactly the same hyperparameters as those used for plates-co-FT found in 9, just applied sequentially twice to first cotraining on the plates dataset, then on the whiteboard. In the continual learning setup, the Merging weight is also always fixed at 50% task-FT, 50% base model.

### v2

@@ 6.5 RETAIN Enables Robust Learning of Multiple Skills Sequentially
Finally, we test whether RETAIN can enable learning multiple skills in sequence, as described in Section˜5.3, and still retain its generalist abilities. We consider learning the two DROID tasks sequentially, first finetuning on plates, and then using this as an initialization to finetune on whiteboard. As outlined in Section˜5.3, RETAIN uses the merged model from the first stage of finetuning as initialization for the second finetuning stage. During evaluation time, we test whether the final policy, after it has sequentially been trained on both tasks, can solve both tasks under ID and OOD evaluations. We compare against co-FT in the sequential learning setting, since it is the strongest baseline at retaining prior knowledge in single-task finetuning (see Fig.˜7). Fig.˜12 shows the performance of the two policies on the two tasks under both ID and OOD settings. We also plot the performance of these two methods in the single-task finetuning setting in dashed lines. These two comparisons serve as the oracle performance ceiling that we expect on these tasks, and is not meant as baselines for the sequential setting. When evaluated on the first task, plates, RETAIN does much better than co-FT in the sequential setting, showing that it is better at retaining its ability to solve the first task even after a second round of finetuning. When evaluated on the second task, whiteboard, RETAIN is also better than co-FT under ID evaluations. RETAIN outperforms co-FT in the sequential setting under all tasks and evaluation types.
@@ 8 Limitations
While we empirically verified that RETAIN works exceptionally well in helping the finetuned model generalize to out-of-distribution variations of the task, we don’t understand the full scope of the reasons why model parameter merging was able to lead to such generalization. This is an interesting area for future work. We have included some discussion in Section˜A.9 of some hypothesess and why previous work found model parameter merging effective for vision and language tasks. Additionally, RETAIN involves an important hyperparameter, the merginge coefficient, that can be tuned. While we find in our real world experiments that RETAIN is robust to different values of this parameter, slight tuning of this parameter is needed. One avenue of future work is determining a good heuristic of how to choose this value.
@@ A.1 DROID Results Details
Fig.˜7 reports performance of two DROID tasks: whiteboard and plates. Here we present the average of the two tasks. The OOD performance shown here is averaged across the val and test scenes.
@@ A.2 LIBERO Results Details
Fig.˜8 reports the average performance of three LIBERO tasks: pot-on-stove, mugs-on-plates, and items-into-box. Here we present the individual performance for all three tasks in Fig.˜14. The OOD performance shown here are averaged only across the two test scene.
@@ A.3 Ablation on Learning Rate and Gradient Steps
Here, we ablate the learning rate and number of gradient steps we take in the task-FT policy in Fig.˜4 to study whether better hyperparameter choices can reduce or resolve overfitting [^1] Following the best practice from Black et al. [8], we always use a learning rate warmup period of $1000$ steps in LIBERO and a consine decay schedule over $30k$ steps to $1/10$ of the peak learning rate. The learning rate we report here is the peak learning rate. . In Fig.˜4, we use learning rate 2.5e52.5e-5. Here in Fig.˜15, we ablate four different learning rates: one greater than the original and two smaller. We evaluate the model at every $100$ gradient steps to also ablate on the number of gradient steps we take. With a larger learning rate, it’s clear that the overfitting issue is more severe, and the performance on all three kinds of evaluations (ID, OOD, Generalist) go down to near zero. With a smaller learning rate, the model still performs well in ID evaluations, and suffers less from forgetting generalist capabilities (measured from Generalist evaluations). This is to be expected because the finetuned model is closer to the pretrained model with a smaller learning rate. However, note that with a smaller learning rate, the OOD evaluation performance is also worse. We plot the OOD performance achieved using the original learning rate in Fig.˜4 as the dotted orange line in Fig.˜15, and the gap between the dotted and solid orange line shows the performance gap in OOD evaluation when we lower the learning rate, possibly due to underfitting. This experiment shows that tuning the learning rate and the number of update steps does not solve the problem of overfitting during finetuning. While lowering the learning rate can retain more generalist knowledge, it does not prevent the gradual loss of it. More importantly, the lower learning rate leads to worse OOD evaluation performance.
@@ A.4 Analysis of Finetuning Path in Parameter Space
To understand why model parameter merging helps, we first try to understand here how the parameters change during fine-tuning. In particular, we are interested in understanding how linear the finetuning path is in parameter space, since model parameter merging only moves weights in linear paths.
To start out, we check the colinearity of the vectors $\theta_{i+1}-\theta_{i}$ and $\theta_{i}-\theta_{i-1}$, where $\theta_{i}$ is the parameter of the finetuned checkpoint at gradient steps 100i100*i (i.e. we plot the difference vector every $100$ steps during finetuning). We measure the colinearity as the cosine similarity between the two difference vectors. Fig.˜16 shows this for the four different learning rates we ablated in Fig.˜15. Since no values are close to $1$, this shows that the changes in parameter space is highly non-linear. This is expected since the parameter trajectory of deep neural networks is often highly non-linear.
@@
Next, we attempt to more directly visualize the path/direction of the parameters during finetuning by projecting them down to 2D with Principal Component Analysis (PCA). Specifically, we again consider the difference of the weight vectors, $X_{i}=\theta_{i+1}-\theta_{i}$, at every $100$ steps during finetuning. In Fig.˜17, we plot the first two principal components of $X_{i}$ in blue, and label the points $i$. Indeed, we see that for all learning rates, the direction of the parameters is highly non-linear. With small learning rates, the direction oscillates a lot; with larger learning rates, the direction bends in a certain direction. In addition, we plot the two principal components of the parameter-merged model as well in orange. As expected, since the the merged model takes a linear path and the finetuned one does not, the two models end in in very different places in the parameter space. This shows that model merging achieves a different solution than any checkpoints on the finetuning path.
Finally, we analyze changes in all directions of the difference vectors $X_{i}$, instead of just the two principal components as shown in Fig.˜17. We take the difference vector matrix $Y=[X_{1};X_{2};...,X_{n}]$ and compute the singular values of $YY^{T}$ and Fig.˜18. If the parameter vectors lie in a linear path, then all difference vectors would point in the same direction and there should only be $1$ non-zero singular value. However, it’s clear from Fig.˜18 that most singular values are non-zero, showing that the path is non-linear in many dimensions. This generalizes the intuition from Fig.˜17 to more dimensions, and shows that model merging indeed achieves a different solution than the finetuning path.
@@
Figure 18: Singular Values of the parameter differences during finetuning. Each subplot corresponds to a different learning rate.
All these analysis experiments go to show that the finetuning path is highly non-linear, and therefore model parameter merging actually results in a different solution than the finetuned models.
@@ A.5.4 Evaluation Details
During evaluations, we additionally binarize the policy’s gripper action to 0/1 (open/close), as well as clip action magnitudes. The policy’s action horizon is 10 environment steps, and during evaluation, we set the open loop horizon to 8: so, during evaluation, we receive 10 actions from the policy, execute the first 8, and then request a new action chunk. We execute the predicted actions at a control frequency of 15 Hz.
@@ A.5.5 Evaluation Criteria: Whiteboard
For all whiteboard evaluations, we use the criteria specified in Table 1 to assign partial success. We perform 10 trials per policy evaluation, for both the ID and OOD evals.
@@ A.5.6 Evaluation Criteria: Plates
For all plates evaluations, we use the criteria specified in Table 2 to assign partial success. We perform 10 trials per policy evaluation, for both the ID and OOD evals.
@@ A.5.7 Generalist Evaluations Details
As detailed in Tables 3, 4, and 5, we measure policies’ generalist capabilities by evaluating them on 44 tasks, distributed throughout 9 distinct scenes and 17 different language instructions. Importantly, to ensure fair comparison, we ensure that the initial conditions, camera angle, lighting, and all other such factors per task are kept the same across the various policies that we evaluate.
Scene Image Task # Trials Randomization Rubric [Uncaptioned image] put the spoon in the dish rack 4 swap spoon and carrot position, 2 evals each 1: pick up spoon; 1.5: move spoon towards dish rack; 2: put spoon in dish rack (anywhere) put carrot in bowl 4 swap spoon and carrot position, 2 evals each 1: pick up carrot; 1.5: move carrot towards bowl; 2: put carrot in bowl put plate in dish rack 2 randomize initial position of the plate in front of the robot 1: pick up plate; 1.5: move plate towards the dish rack; 2: put plate into dish rack (anywhere) [Uncaptioned image] wipe the table 2 cloth initially on the left and right side of the open area 1: move down towards cloth; 2: perform lateral ""wiping-style"" motion [Uncaptioned image] put the plate on the table 2 plate initially on different dish rack holders (middle and end) 1: moves towards red plate; 1.5: picks up plate; 2: places / drops plate onto the table [Uncaptioned image] clean up the table 2 randomize initial position of paper ball on table 1: picks up paper ball; 1.5 moves paper ball towards brown bin; 2: puts paper ball into bin
Table 3: DROID Generalist evaluation tasks grouped by scene (Scenes 1-4). Each task contains associated trial counts, randomization details, and evaluation rubrics.
Scene Image Task # Trials Randomization Rubric [Uncaptioned image] close the drawer 4 two top, two bottom drawer 1: moves towards the drawer; 2: closes drawer [Uncaptioned image] put the stapler on the notebook 2 put stapler in higher and lower position on the table 1: picks up stapler; 2: puts stapler on notebook put stapler in the drawer 4 put stapler in higher and lower position on the table, open top and bottom drawer 1: picks up the stapler; 2: puts stapler into the drawer [Uncaptioned image] clean the whiteboard 2 initial eraser position on the left and right of whiteboard 1: pick up eraser; 2: perform wiping motions on the whiteboard; 3: erase the full smiley [Uncaptioned image] put the marker in the cup 4 swap initial position of marker and cup, two local modifications each 1: picks up marker; 1.5: rotates arm to put marker in roughly upright position; 2: puts marker in cup
Table 4: DROID Generalist evaluation tasks grouped by scene (Scenes 5-8). Each task contains associated trial counts, randomization details, and evaluation rubrics.
@@
Table 5: DROID Generalist evaluation tasks grouped by scene (Scene 9). Each task contains associated trial counts, randomization details, and evaluation rubrics.
@@ A.6.1 LIBERO Pretraining
In order to obtain a base-model to serve as the starting point for RETAIN in LIBERO, we pretrain $\pi_{0}$ on a mixture of LIBERO datasets. Specifically, we use 90 tasks from LIBERO-90, 9 tasks from LIBERO-object, 9 tasks from LIBERO-spatial, and 9 tasks from LIBERO-goal, for a total of 117 tasks in our pretraining dataset. For all LIBERO datasets, pretraining and finetuning, we utilize pre-processed RLDS versions gathered from here and here. These datasets consist of LIBERO demonstrations that have been preprocessed to upscale images, filter out transitions with idle actions, and remove failure trajectories. More details, such as hyperparameter choices, for the pretraining stage itself can be found in A.7. All subsequent fine-tuning and evaluation uses the normalization stats of this pretraining dataset, applying mean/standard-deviation normalization.
@@ A.6.2 LIBERO Finetuning Datasets
As mentioned in 6.1, we finetune on 3 tasks from LIBERO-10: pot-on-stove, mugs-on-plates, and items-into-basket. For each dataset, we obtain pre-processed and filtered versions as described above. 19 highlights what these finetuning tasks, and thus also our ID evals, look like.
@@ A.6.3 LIBERO Evaluation Details
We use the LIBERO simulator for evaluation. Each ID eval is conducted for 20 episodes. Each OOD eval is conducted with 5 seeds, 10 episodes/seed. 20, 21, and 22 show the 3 types of OOD variations we test for in each of the 3 tasks.
The Generalist evals consists of 20 tasks, 5 each from LIBERO-object, LIBERO-spatial, LIBERO-goal, and LIBERO-90, and each task is tested for 10 episodes. Table 6 highlights a few tasks per LIBERO eval suite that we use in our generalist evals.
@@ A.7.1 Choosing Merging Coefficient $\alpha$
For each task that we consider (e.g. mugs-on-plate in LIBERO, plates in DROID), we test policy performance on several scenes, which are different variations of the same task with different objects, distractors, and backgrounds etc. To choose what values of $\alpha$ we use, we use one OOD scene as the “validation” scene, and tune the hyperparameter $\alpha$ for best performance on that validation scene. Then, we use the rest of the OOD scenes as the “test” scenes, and report the performance of all methods only on the test scenes. In DROID experiments, we only tune $\alpha\in\{0.25,0.5,0.75\}$ on the validation scene, while in LIBERO experiments we tune $\alpha\in\{0.9,0.8,0.7,0.6,0.5,0.4,0.3,0.2,0.1\}$ since it is cheap to do so. We find that on DROID $\alpha=0.5$ typically tends to perform well across tasks.
@@ A.7.3 Real-world Experiments’ Hyperparameters
Finally, in our continual learning experiment, we use exactly the same hyperparameters as those used for plates-co-FT found in 10, just applied sequentially twice to first cotraining on the plates dataset, then on the whiteboard. In the continual learning setup, the Merging weight is also always fixed at 50% task-FT, 50% base model.